# DepthART: Monocular Depth Estimation as Autoregressive Refinement Task

### v2

@@ Abstract
Despite recent success in discriminative approaches in monocular depth estimation its quality remains limited by training datasets. Generative approaches mitigate this issue by leveraging strong priors derived from training on internet-scale datasets. Recent studies have demonstrated that large text-to-image diffusion models achieve state-of-the-art results in depth estimation when fine-tuned on small depth datasets. Concurrently, autoregressive generative approaches, such as the Visual AutoRegressive modeling (VAR), have shown promising results in conditioned image synthesis. Following the visual autoregressive modeling paradigm, we introduce the first autoregressive depth estimation model based on the visual autoregressive transformer. Our primary contribution is DepthART – a novel training method formulated as Depth Autoregressive Refinement Task. Unlike the original VAR training procedure, which employs static targets, our method utilizes a dynamic target formulation that enables model self-refinement and incorporates multi-modal guidance during training. Specifically, we use model predictions as inputs instead of ground truth token maps during training, framing the objective as residual minimization. Our experiments demonstrate that the proposed training approach significantly outperforms visual autoregressive modeling via next-scale prediction in the depth estimation task. The Visual Autoregressive Transformer trained with our approach on Hypersim achieves superior results on a set of unseen benchmarks compared to other generative and discriminative baselines.
Figure 1: We present the Depth Autoregressive Transformer for monocular depth estimation, trained using our novel procedure formulated as the Depth Autoregressive Refinement Task – DepthART. Our model iteratively enhances the depth map by predicting next-scale residuals, resulting in a highly detailed final estimate.
Introduction
@@
Monocular depth estimation (MDE) is a fundamental problem in computer vision. Depth maps provide a compact intermediate scene representation useful for decision making in physical surroundings. Recovering depth data from a single image promises a high practical value for different applications including spatial vision intelligence (Wang et al. 2019; Godard et al. 2019), autonomous driving (Wang et al. 2019; Godard et al. 2019) and robotics (Wofk et al. 2019).
@@
Early learning-based approaches (Eigen, Puhrsch, and Fergus 2014) tackle the monocular depth estimation problem as a supervised regression task. However, these methods were domain-specific (Silberman et al. 2012; Geiger, Lenz, and Urtasun 2012) and heavily relied on annotated datasets. As a result, they were subject to a limited generalization ability caused by a low amount of annotated data available. Recent techniques suggested different tricks challenging this limitation. MiDaS  (Ranftl et al. 2020) proposed to mitigate this issue by using an affine invariant depth training scheme on a mixture of datasets. While newer approaches proposing annotated data sources still appear (Yang et al. 2024), the acquisition of accurate depth annotations at scale remains challenging.
@@
Recent studies (Ke et al. 2024; Fu et al. 2024) have highlighted the effectiveness of text-to-image diffusion models, originally trained on internet-scale image-caption datasets, as priors for monocular depth estimation. These approaches involve fine-tuning a pretrained diffusion model on a smaller, synthetic dataset with depth annotations, resulting in models that generate accurate and highly detailed depth maps. Concurrently, advancements in autoregressive models, such as the Visual AutoRegressive modeling (VAR) (Tian et al. 2024) and LLaMA-Gen (Sun et al. 2024), have demonstrated the capability of these models to generate high-quality images in class- or text-guided settings. These findings motivate an exploration of autoregressive generative techniques for depth estimation, offering a promising new direction.
@@
In this work, we introduce a novel approach to monocular depth estimation based on the Visual AutoRegressive modeling (Tian et al. 2024). Our core contribution is the novel training procedure formulated as Depth Autoregressive Refinement Task. Our approach constructs dynamic targets using the model’s own predictions, rather than relying on ground truth token maps during training. By framing the objective as residual minimization and using model predictions as inputs, we bridge the gap between training and inference stages in autoregressive modeling, leading to enhanced depth estimation quality. We validate our model extensively comparing it with popular baselines under similar conditions. To the best of our knowledge, this is the first autoregressive depth estimation model. Moreover, it performs on-par or superior compared with popular depth estimation baselines.
@@
    We introduce a novel application of autoregressive image modeling for depth estimation by developing the depth autoregressive transformer.
@@
    We propose a new training paradigm for depth estimation, termed the Depth Autoregressive Refinement Task (DepthART), which facilitates self-refinement and incorporates multi-modal guidance during training.
@@
    We demonstrate, through extensive experiments, that the depth autoregressive transformer trained with DepthART achieves competitive or superior performance compared to existing baselines across several benchmarks not seen during training.
Related work
Monocular depth estimation
@@
Learning-based monocular depth estimation approaches can be broadly categorized into two main branches: metric and relative depth estimation methods. Metric depth estimation (Laina et al. 2016; Alhashim and Wonka 2018; Bhat, Alhashim, and Wonka 2021, 2022; Yin et al. 2023) focuses on regressing absolute predictions at a metric scale. These models are typically trained on small, domain-specific datasets, which limits their ability to generalize efficiently across diverse environments. At the same time, relative depth estimation methods aim to estimate depth up to unknown shift and scale (SSI) or just unknown scale (SI). MiDaS  (Ranftl et al. 2020) introduced shift and scale invariant depth training on a mixture of several domain-specific datasets, significantly improving model generalization. Despite it, the depth predictions remained geometrically incomplete, i.e. point clouds cannot be built using model predictions. GP2  (Patakin et al. 2022) addressed this limitation by proposing an end-to-end training scheme that estimates a scale-invariant, geometry-preserving depth maps. Meanwhile, two-stage pipelines were developed to reduce shift ambiguity in depth maps at the second stage (Yin et al. 2021) or to upgrade the depth map to metric scale (Bhat et al. 2023). Further advancements in the field have been driven by the integration of various priors (Patil et al. 2022; Yang et al. 2024; Yin et al. 2023), improvements in architectural designs (Ranftl, Bochkovskiy, and Koltun 2021; Agarwal and Arora 2023; Ning and Gan 2023) and the expansion of training data (Yang et al. 2024).
Generative modeling
@@
Recently, diffusion models have demonstrated their versatility across various computer vision tasks, including image generation  (Ho, Jain, and Abbeel 2020; Rombach et al. 2022; Ho et al. 2022), video generation  (Blattmann et al. 2023a; Bar-Tal et al. 2024; Blattmann et al. 2023b), or 3D objects modeling  (Poole et al. 2022; Wang et al. 2023b; Lin et al. 2023; Melas-Kyriazi et al. 2023). Beyond these applications, diffusion models have also been successfully employed in other problems, such as depth estimation (Saxena et al. 2023; Duan, Guo, and Zhu 2023; Saxena et al. 2024), image segmentation (Wang et al. 2023a; Amit et al. 2021) and object detection (Chen et al. 2023). Notably, Marigold  (Ke et al. 2024) and GeoWizard  (Fu et al. 2024) have demonstrated that the Stable Diffusion model (Rombach et al. 2022), pretrained on the large-scale image-caption dataset LAION-5B (Schuhmann et al. 2022), can produce high-quality depth maps after minor finetuning. This highlights the potential of utilizing pretrained generative models to enhance depth estimation accuracy and robustness across different domains.
Autoregressive modeling
@@
While diffusion models remain one of the most widely-used generative approach, recent advancements in autoregressive models have shown significant promise for various generative tasks (Yu et al. 2023; Sun et al. 2024; Tian et al. 2024; Ma et al. 2024). These methods rely on discrete token-based image representations typically generated by VQ-VAE (Van Den Oord, Vinyals et al. 2017) or its derivatives. These derivatives often include architectural enhancements (Yu et al. 2021; Razavi, Van den Oord, and Vinyals 2019), additional masking techniques (Huang et al. 2023), or the incorporation of adversarial and perceptual losses (Esser, Rombach, and Ommer 2021). Autoregressive image synthesis is generally formulated as the sequential generation of tokens (Esser, Rombach, and Ommer 2021), followed by decoding them from the VQ space. Many approaches employ the GPT-2 (Radford et al. 2019) decoder-only architecture to predict sequences of VQ-VAE tokens (Esser, Rombach, and Ommer 2021; Chang et al. 2022; Ramesh et al. 2021; Chang et al. 2023; Yu et al. 2022). However recent works (Tian et al. 2024; Ma et al. 2024) have introduced the concept of predicting multi-scale token maps rather than token sequences. This approach reduces the risk of structural degradation and decreases the generation time for high-resolution images, enabling high-quality class- and text-conditioned image generation.
Preliminary
@@
Figure 2: VQ-VAE (Tian et al. 2024) provides a single decomposition, while there are different trajectories resulting the same features. In contrast, DepthART enables multi-modal training, offering diverse refinement paths compared to the unimodal guidance used in the Visual AutoRegressive modeling (VAR).
@@
Figure 3: We highlight the key differences between the original VAR approach (left) and our proposed training approach DepthART (right). In the VAR approach, quantized token maps provided by VQ-VAE serve as both inputs and targets during training. Our DepthART method introduces a refinement process (highlighted in the red box), where the model self-refines by using its predicted token maps as inputs instead of predefined VQ-VAE scales. The targets are defined as the quantized residuals between the encoded depth features $f_{D}$ and the cumulative model predictions up to the current scale. Depth features $f_{D}$ are extracted from the VQ-VAE encoder without undergoing quantization.
@@
Typically autoregressive image generation involves predicting image tokens in a raster scan order. However, recent work (Tian et al. 2024) introduced a novel autoregressive training approach for class-conditional image generation - visual autoregressive modeling. Instead of predicting tokens individually, they proposed generating token maps with varying scale. Each predicted token map progressively increases in resolution compared to the previous one, resulting in a scale-wise decomposition of the image. Specifically, an image ${\bm{I}}$ is modelled as a sequence of $\{x_{1},x_{2},\dots,x_{K}\}$, where $x_{k}$ is a token map of size $s_{k}=(h_{k},w_{k})$. The visual autoregressive transformer based on GPT-2 is then trained to maximize a class-conditioned likelihood:
@@
Effective autoregressive image modeling relies on discretizing images into a finite set of tokens, a task closely related to image compression. Image compression methods have evolved from linear projections to advanced neural approaches, such as vector quantized variational autoencoders (VQ-VAE) (Van Den Oord, Vinyals et al. 2017). VQ-VAE approach combines an encoder-decoder neural network with a learnable token codebook of finite size. The encoder compresses the input image into a latent space, which is then quantized by mapping the latent vectors to the closest entries in the codebook. This quantized representation is then decoded back into the image space, with the entire process trained end-to-end using a log-likelihood objective over the image distribution.
Visual autoregressive model training requires multi-scale image decomposition represented as a sequence of token maps. Differently from original VQ-VAE (Van Den Oord, Vinyals et al. 2017) an input image ${\bm{I}}$ is quantized by encoder ${\mathcal{E}}$ into $K$ token maps $\{x_{1},x_{2},\dots,x_{K}\}$ with resolutions $\{(h_{k},w_{k})\}_{k=1,K}$. Quantization operation $Q[\cdot]$ is performed using the same codebook $Z$ regardless of scale. Combined using upsample-convolution operators $\eta_{i}$ token maps are assumed to sum up into continuous features ${\mathcal{E}}({\bm{I}})$. Accordingly, the $k$-th scale map is calculated as a scaled and quantized residual of the extracted features:
$r_{k}={\mathcal{E}}({\bm{I}})-\sum_{i=1}^{k-1}\eta_{i}(x_{i})$ (2)
@@
Eventually, decoder ${\mathcal{D}}$ recovers reconstructed image $\hat{{\bm{I}}}$ from given token maps:
$\hat{{\bm{I}}}={\mathcal{D}}\bigg{(}\sum_{k=1}^{K}\eta_{k}(x_{k})\bigg{)}$ (4)
In our approach, we adapt a pretrained modification of VQ-VAE tailored specifically for the visual autoregressive modeling transformer (VAR) (Tian et al. 2024). While originally designed for colored images, we observe that VQ-VAE (Tian et al. 2024) can be applied to encode depth maps as well (see  fig. 1).
Method
In this work, we formulate the monocular depth estimation task as an image-conditioned autoregressive generation problem. Inspired by VAR, we develop the depth autoregressive transformer that generates depth token maps given image conditioning. As our primary contribution, we introduce the novel training procedure formulated as Depth Autoregressive Refinement Task – DepthART.
Figure 4: Depth autoregressive transformer trained with DepthART demonstrates superior reconstruction quality at every scale (left) compared to the standart VAR training procedure, achieving a 1.7x improvement in reconstruction quality. Since we do not finetune the VQ-VAE, its end-to-end reconstruction quality is shown as a soft limitation on reconstruction error. Additionally, the overall predicted token probability distributions exhibit higher average entropy (right), indicating that our training procedure promotes multi-modality.
Depth Autoregressive Refinement Task.
The original Visual AutoRegressive modeling (Tian et al. 2024) relies on scale-wise decomposition of image provided by pre-trained VQ-VAE encoder. During training the model predicts next-scale token map from the previous sequence of ground truth token maps. The guidance objective is a cross-entropy loss between VAR prediction and the same scale token map provided by VQ-VAE. Inputting ground-truth token maps results into discrepancies between training and inference process, and accumulation of errors during inference.
We address this issue and reformulate training objective as a Depth Autoregressive Refinement Task (DepthART). Our main goal is to enable model self-refinement during training. Hence, we construct inputs and targets dynamically from model predictions. Let’s consider an input image ${\bm{I}}$ with corresponding ground truth-depth map ${\bm{D}}$. We firstly encode an image into a series of token maps $\{x_{1},x_{2},\dots x_{K}\}$ provided by VQ-VAE (Tian et al. 2024). Resulting image token maps are fed to the model input as a starting sequence and serve as a conditioning for depth map estimation. Constructing dynamic supervision targets in our approach starts with performing model inference for given image token maps. We denote predicted depth token maps as $\{z_{1},z_{2},\dots,z_{K}\}$:

### v3

@@ Abstract
Monocular depth estimation has seen significant advances through discriminative approaches, yet their performance remains constrained by the limitations of training datasets. While generative approaches have addressed this challenge by leveraging priors from internet-scale datasets, with recent studies showing state-of-the-art results using fine-tuned text-to-image diffusion models, there is still room for improvement. Notably, autoregressive generative approaches, particularly Visual AutoRegressive modeling, have demonstrated superior results compared to diffusion models in conditioned image synthesis, while offering faster inference times. In this work, we apply Visual Autoregressive Transformer (VAR) to the monocular depth estimation problem. However, the conventional GPT-2-style training procedure (teacher forcing) inherited by VAR yields suboptimal results for depth estimation. To address this limitation, we introduce DepthART - a novel training method formulated as a Depth Autoregressive Refinement Task. Unlike traditional VAR training with static inputs and targets, our method implements a dynamic target formulation based on model outputs, enabling self-refinement. By utilizing the model’s own predictions as inputs instead of ground truth token maps during training, we frame the objective as residual minimization, effectively reducing the discrepancy between training and inference procedures. Our experimental results demonstrate that the proposed training approach significantly enhances the performance of VAR in depth estimation tasks. When trained on Hypersim dataset using our approach, the model achieves superior results across multiple unseen benchmarks compared to existing generative and discriminative baselines.
1 Introduction
@@
Monocular depth estimation (MDE) is a fundamental problem in computer vision. Depth maps provide a compact intermediate scene representation useful for decision making in physical surroundings. Recovering depth data from a single image promises a high practical value for different applications including spatial vision intelligence Wang et al. (2019), autonomous driving Wang et al. (2019) and robotics Wofk et al. (2019).
@@
Early learning-based approaches Eigen et al. (2014) tackle the monocular depth estimation problem as a supervised regression task. However, these methods were domain-specific Silberman et al. (2012) and heavily relied on annotated datasets. As a result, they were subject to a limited generalization ability caused by a low amount of annotated data available. Recent techniques suggested different tricks challenging this limitation. MiDaS  Ranftl et al. (2020) proposed to mitigate this issue by using an affine invariant depth training scheme on a mixture of datasets. While newer approaches proposing annotated data sources still appear Yang et al. (2024), the acquisition of accurate depth annotations at scale remains challenging.
@@
Recent studies Ke et al. (2024); Fu et al. (2024) have highlighted the effectiveness of text-to-image diffusion models, originally trained on internet-scale image-caption datasets, as priors for monocular depth estimation. These approaches involve fine-tuning a pretrained diffusion model on a smaller, synthetic dataset with depth annotations, resulting in models that generate accurate and highly detailed depth maps. Concurrently, advancements in autoregressive models, such as the Visual AutoRegressive modeling  Tian et al. (2024) and LLaMA-Gen Sun et al. (2024), have demonstrated the capability of these models to generate high-quality images in class- or text-guided settings. These findings motivate an exploration of autoregressive generative techniques for depth estimation, offering a promising new direction.
@@
In this work, we introduce [^1] https://bulatko.github.io/depthart-pp/ a novel approach to monocular depth estimation based on the Visual AutoRegressive modeling Tian et al. (2024). Our core contribution is the novel training procedure formulated as Depth Autoregressive Refinement Task. Our approach constructs dynamic targets using the model’s own predictions, rather than relying on ground truth token maps during training. By framing the objective as residual minimization and using model predictions as inputs, we bridge the gap between training and inference stages in autoregressive modeling, leading to enhanced depth estimation quality. We validate our model extensively comparing it with popular baselines under similar conditions. To the best of our knowledge, this is the first autoregressive depth estimation model. Moreover, it performs on-par or superior compared with popular depth estimation baselines.
@@
    We introduce a novel application of Visual Autoregressive Transformer for depth estimation.
@@
    We propose a new training paradigm for depth estimation, termed the Depth Autoregressive Refinement Task (DepthART), which facilitates self-refinement and bridges the gap between training and inference procedures.
@@
    We demonstrate, through extensive experiments, that the Visual Autoregressive Transformer trained with DepthART achieves competitive or superior performance compared to existing baselines across several benchmarks not seen during training.
2 Related Work
2.1 Monocular Depth Estimation
@@
Learning-based monocular depth estimation approaches can be broadly categorized into two main branches: metric and relative depth estimation methods. Metric depth estimation Laina et al. (2016); Bhat et al. (2021); Yin et al. (2023) focuses on regressing absolute predictions at a metric scale. These models are typically trained on small, domain-specific datasets, which limits their ability to generalize efficiently across diverse environments. At the same time, relative depth estimation methods aim to estimate depth up to unknown shift and scale (SSI) or just unknown scale (SI). MiDaS  Ranftl et al. (2020) introduced shift and scale invariant depth training on a mixture of several domain-specific datasets, significantly improving model generalization. Despite it, the depth predictions remained geometrically incomplete, i.e. point clouds cannot be built using model predictions. GP2  Patakin et al. (2022) addressed this limitation by proposing an end-to-end training scheme that estimates a scale-invariant, geometry-preserving depth maps. Meanwhile, two-stage pipelines were developed to reduce shift ambiguity in depth maps at the second stage Yin et al. (2021) or to upgrade the depth map to metric scale Bhat et al. (2023). Further advancements in the field have been driven by the integration of various priors Yang et al. (2024); Yin et al. (2023), improvements in architectural designs Ranftl et al. (2021); Agarwal and Arora (2023); Ning and Gan (2023) and the expansion of training data Yang et al. (2024).
2.2 Generative Modeling
@@
Recently, diffusion models have demonstrated their versatility across various computer vision tasks, including image generation  Ho et al. (2020); Rombach et al. (2022), video generation  Blattmann et al. (2023), or 3D objects modeling  Melas-Kyriazi et al. (2023). Beyond these applications, diffusion models have also been successfully employed in other problems, such as depth estimation Duan et al. (2023), image segmentation Wang et al. (2023) and object detection Chen et al. (2023). Notably, Marigold  Ke et al. (2024) and GeoWizard  Fu et al. (2024) have demonstrated that the Stable Diffusion model Rombach et al. (2022), pretrained on the large-scale image-caption dataset LAION-5B Schuhmann et al. (2022), can produce high-quality depth maps after minor finetuning. This highlights the potential of utilizing pretrained generative models to enhance depth estimation accuracy and robustness across different domains.
2.3 Autoregressive Modeling
@@
While diffusion models remain one of the most widely-used generative approach, recent advancements in autoregressive models have shown significant promise for various generative tasks Yu et al. (2023); Sun et al. (2024); Tian et al. (2024). These methods rely on discrete token-based image representations typically generated by VQ-VAE Van Den Oord et al. (2017) or its derivatives. These derivatives often include architectural enhancements Yu et al. (2021), additional masking techniques Huang et al. (2023), or the incorporation of adversarial and perceptual losses Esser et al. (2021). Autoregressive image synthesis is generally formulated as the sequential generation of tokens Esser et al. (2021), followed by decoding them from the VQ space. Many approaches employ the GPT-2 Radford et al. (2019) decoder-only architecture to predict sequences of VQ-VAE tokens Esser et al. (2021); Chang et al. (2023). However recent works Tian et al. (2024) have introduced the concept of predicting multi-scale token maps rather than token sequences. This approach reduces the risk of structural degradation and decreases the generation time for high-resolution images, enabling high-quality class- and text-conditioned image generation.
3 Preliminary
@@
Figure 1: VQ-VAE provides a single decomposition, while there are different trajectories that can result in the same features. The usage of teacher forcing during training leads to selecting only one trajectory, given by the VQ-VAE decomposition. In contrast, DepthART enables the use of other possible decompositions, offering diverse refinement paths compared to the unique guidance used in teacher forcing.
@@
Figure 2: DepthART (right). We also demonstrate how our training procedure better aligns with the inference process (shown at the bottom), as in both training and inference, the model operates on its own predictions as inputs. In the teacher forcing approach, quantized token maps provided by VQ-VAE serve as both inputs and targets during training. Our DepthART method introduces a refinement process (highlighted in the red box), where the model self-refines by using its predicted token maps as inputs instead of predefined VQ-VAE scales. The targets are defined as the quantized residuals between the encoded depth features $f_{D}$ and the cumulative model predictions up to the current scale. Depth features $f_{D}$ are extracted from the VQ-VAE encoder without undergoing quantization.
@@
Typically autoregressive image generation involves predicting image tokens in a raster scan order. However, recent work Tian et al. (2024) introduced a novel autoregressive training approach for class-conditional image generation - visual autoregressive modeling. Instead of predicting tokens individually, they proposed generating token maps with varying scale. Each predicted token map progressively increases in resolution compared to the previous one, resulting in a scale-wise decomposition of the image. Specifically, an image I is modelled as a sequence of $\{x_{1},x_{2},\dots,x_{K}\}$, where $x_{k}$ is a token map of size $s_{k}=(h_{k},w_{k})$. The visual autoregressive transformer based on GPT-2 is then trained to maximize a class-conditioned likelihood:
@@
Effective autoregressive image modeling relies on discretizing images into a finite set of tokens, a task closely related to image compression. Image compression methods have evolved from linear projections to advanced neural approaches, such as vector quantized variational autoencoders (VQ-VAE) Van Den Oord et al. (2017). VQ-VAE approach combines an encoder-decoder neural network with a learnable token codebook of finite size. The encoder compresses the input image into a latent space, which is then quantized by mapping the latent vectors to the closest entries in the codebook. This quantized representation is then decoded back into the image space, with the entire process trained end-to-end using a log-likelihood objective over the image distribution.
Visual autoregressive model training requires multi-scale image decomposition represented as a sequence of token maps. Differently from original VQ-VAE Van Den Oord et al. (2017) an input image I is quantized by encoder ${\mathcal{E}}$ into $K$ token maps $\{x_{1},x_{2},\dots,x_{K}\}$ with resolutions $\{(h_{k},w_{k})\}_{k=1,K}$. Quantization operation $Q[\cdot]$ is performed using the same codebook $Z$ regardless of scale. Combined using upsample-convolution operators $\eta_{i}$ token maps are assumed to sum up into continuous features ${\mathcal{E}}({\textit{I}})$. Accordingly, the $k$-th scale map is calculated as a scaled and quantized residual of the extracted features:
$r_{k}={\mathcal{E}}({\textit{I}})-\sum_{i=1}^{k-1}\eta_{i}(x_{i})$ (2)
@@
${\mathcal{S}}$ scaling operation.
Eventually, decoder ${\mathcal{D}}$ recovers reconstructed image $\hat{{\textit{I}}}$ from given token maps:
$\hat{{\textit{I}}}={\mathcal{D}}\bigg{(}\sum_{k=1}^{K}\eta_{k}(x_{k})\bigg{)}$ (4)
In our approach, we adapt a pretrained modification of VQ-VAE tailored specifically for the visual autoregressive modeling transformer (VAR) Tian et al. (2024). While originally designed for colored images, we observe that VQ-VAE Tian et al. (2024) can be applied to encode depth maps as well.
4 Method
In this work, we formulate the monocular depth estimation task as an image-conditioned autoregressive generation problem. We leverage the Visual Autoregressive Transformer (VAR) for monocular depth estimation, adapting it to generate depth token maps from input images. As our primary contribution, we introduce the novel training procedure formulated as Depth Autoregressive Refinement Task – DepthART.
Figure 3: VAR trained with DepthART demonstrates superior reconstruction quality at every scale compared to the teacher forcing training procedure, achieving a 1.7x improvement in reconstruction quality. Since we do not finetune the VQ-VAE, its end-to-end reconstruction quality is shown as a soft limitation on reconstruction error.
4.0.1 Depth Autoregressive Refinement Task
The original Visual AutoRegressive modeling Tian et al. (2024) relies on scale-wise decomposition of image provided by pre-trained VQ-VAE encoder. During training the model predicts next-scale token map from the previous sequence of ground truth token maps. Ground truth token maps are computed using VQ-VAE (Eq. 3, 2). The guidance objective is a cross-entropy loss between VAR prediction and the same scale token map provided by VQ-VAE. Inputting ground-truth token maps during training creates a discrepancy with inference, as the model never learns to account for its own prediction errors from previous steps. However, since the final feature map is constructed as a sum of all scale contributions, and this sum can be achieved through different combinations of addends, we can reformulate the problem: at each step, we can compute the difference between the current accumulated sum and the target feature map, allowing us to define the target for each scale as a residual that needs to be predicted to reach the final feature value.
We address this issue and reformulate training objective as a Depth Autoregressive Refinement Task (DepthART). Our main goal is to enable model self-refinement during training. Hence, we construct inputs and targets dynamically from model predictions. Let’s consider an input image I with corresponding ground truth-depth map D. We firstly encode an image into a series of token maps $\{x_{1},x_{2},\dots x_{K}\}$ Eq. 3 provided by VQ-VAE Tian et al. (2024). Resulting image token maps are fed to the model input as a starting sequence and serve as a conditioning for depth map estimation. Constructing dynamic supervision targets in our approach starts with performing model inference for given image token maps. We denote predicted depth token maps as $\{z_{1},z_{2},\dots,z_{K}\}$: